OpenStudy (anonymous):

what is the inverse of this equation A(b)=10*(b=7)/2 answers: A. a/7-5 B. a/7+5 C. a/5-7 D. a/5+7

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

is it: \(A(b)=10(b-7)/2\) ?

OpenStudy (anonymous):

yeah sorry for the typo

OpenStudy (anonymous):

then express b in terms of a: \(a=10(b-7)/2\) \(2a=10(b-7)\) \(\frac{1}{5}a=b-7\) \(\frac{a}{5}+7=b\)

OpenStudy (anonymous):

ok i think i understand...